the evidence of an illegal gun operation consisted of two different types of guns. the pistols are worth $565 a piece and the rifles are worth $795 a piece. If there are two times as many rifles as there are pistols, and all the guns combined are worth $75,425, how many rifles and pistols are being held in evidence?
Answer by addingup(3677)

R=2P
565P+795(2P)=75425
565P+1590P=75425
2155P=75425
P=75425/2155 35 and R=2(35)=70
^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^
John